• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 16:07:14 on March22,2005

    Entry number 142196

    keyword=shift summary - day

    2870 ended 08:39 200K ev
    08:50 HRS-R angle is updated (was 0 due to previous epics crash)
    2871 ended 09:29 200K ev
    09:31 MCC will try to increase current to 40 uA (previously about 35)
    2872 ended 10:13 200K ev
    2873 ended 11:00 200K ev
    lumi 3 HV turned on to -1100 V (by Riad for his tests) should not affect anything
    writing data to a new computer disk. rebooting daq computer to do this. ROM will halog it.
    11:30 lumi 3 HV off
    2874 ended 11:55 200k events
    2875 ended 12:43 200k events
    2876 ended 13:28 200k events
    2877 ended 14:18 200k events
    2878 ended 15:02 200k events
    2879 ended 15:12 due to lost beam at 10k events
    15:21 start run 2880